# A Comprehensive Study on Large Language Models for Mutation Testing

### v4

@@ 4.3. Effectiveness on High-Order Mutation Generation
Previously, we only evaluated LLMs in generating first-order mutations (FOM). To explore the effectiveness of LLMs in generating high-order mutations (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating the recent mainstream models, including GPT-3.5, GPT-4o, and DeepSeek-V3-761b.
@@
We modified the original prompt in Figure 1 to explicitly request HOM generation. Table 20 shows the results. For the effectiveness metrics, HOMs generally achieve average Ochiai scores and higher coupling rates across these models, indicating closer semantic alignment with real bugs. However, FOMs achieve higher real bug detectability on all models. For the validity metrics, HOMs exhibit lower duplicate mutation rates, suggesting improved mutation diversity. However, FOMs exhibit significantly higher compilability rates, which aligns with the expectation that larger code changes are more prone to introducing compilation errors. Our preliminary experiments suggest that strengthening LLM-based HOM generation is promising, as HOMs are semantically closer to real bugs and face more challenges in maintaining syntactic correctness.
@@ 4.4. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches and suggest potential improvements.
@@ 4.4.1. Advantages of LLM-based Approaches
Our study demonstrates that LLMs can generate high-quality mutations for mutation testing at an acceptable cost. The key advantage of LLMs lies in their ability to produce mutations with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
Additionally, LLMs could generate syntactically diverse mutations compared with traditional approaches, which is potentially the reason for generating high-coupling mutations. For example, Listing 1 presents two compilable mutations generated by GPT-3.5 for Chart-1, which are theoretically beyond the capabilities of traditional approaches. In Mutation-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly. This mutation is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement. Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutations. In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
Additionally, our study explores how different prompt designs, prompt turbulence, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
Moreover, our study demonstrates that LLMs have clear potential in generating HOMs. Compared with first-order mutations (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability. The results highlight that strengthening LLM-based HOM generation is therefore a promising research direction.
@@ 4.4.2. Advantages of Traditional Approaches
Despite the advantages of LLM-generated mutations, traditional approaches still exhibit strengths in reliability and efficiency. They produce fewer non-compilable and duplicate mutations, with PIT achieving a 100% compilability rate and a 0% duplicate mutation rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutation rate. In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutation rates. Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutation rate, further avoiding redundant costs. This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
@@ 4.4.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate that there is still substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
Moreover, exploring the applications of LLM-generated mutation in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL). Our study has shown that LLM-generated mutations often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
Another important direction is cross-language generalization. While our study primarily focuses on Java, mutation testing is equally critical in languages such as C/C++, Python, JavaScript, and Rust. Traditional mutation testing tools usually require language-specific implementations, leading to duplicated engineering efforts and limited portability. In contrast, modern LLMs possess inherent multilingual capabilities, as they are pre-trained on large-scale, cross-language code corpora. This enables them to naturally transfer mutation strategies across programming languages without retraining from scratch. Future work should investigate how well LLMs can exploit this cross-language knowledge, and how prompts, training data, or filtering strategies can be adapted to different language paradigms (e.g., strongly typed vs. dynamically typed). Advancing towards multilingual mutation generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
@@ 4.5. Limitations of Our Study
First, while we evaluated a comprehensive set of static and behavioral metrics, we did not investigate more advanced dynamic indicators. For instance, the phenomenon of subsuming mutations (Garg et al., 2022; Diniz et al., 2021), where one mutation is always killed by any test that kills another, has been studied in the mutation testing literature as a way to assess better redundancy and the true fault-detection power of test suites. Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutations beyond the metrics reported in this paper. Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
Second, our experiments are limited to the Java programming language. Although Java is the most widely studied language in mutation testing, it does not cover the diversity of programming paradigms. Mutation behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments. Future research should examine the generalizability of LLM-based mutation generation across multiple languages. Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutation generation a promising direction for broadening the applicability of our findings.
Third, our study relies on fixed few-shot examples, which restricts LLMs to producing context-specific mutations for the given code snippets. In addition, we impose no restrictions on the mutation operators applied by LLMs, whereas rule-based approaches rely on predefined and relatively simple operators. Consequently, it is still an open question how LLMs would perform if constrained to the same operator set as rule-based approaches.
@@ 5. Threats to Validity
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek). Moreover, our study further adopts two state-of-the-art LLM-based mutation generation approaches (i.e., BugFarm and LLMorpheus) as baselines.
@@
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014b) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of the most LLMs we use (expect DeepSeek-V3-761b only) and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek-V3-761b (72), DeepSeek-V2-236b (66), BugFarm (59), LLMorpheus (26), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek-V3-761b (35), DeepSeek-V2-236b (32), BugFarm (22), LLMorpheus (19), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with -value less than 0.05, indicating their performance is similar on the two datasets.
@@
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompt design, prompt turbulence, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs.
@@
The subjective nature of human decisions in labeling for equivalent mutations, non-compilation errors, and alive mutations is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
@@ 6. Related Work
6.1. Mutation Testing and Mutation Generation
@@
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978), one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
@@
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to be coupled with real faults/bugs to allow these techniques to be effective.
@@
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
@@
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. LEAM++, the extension of LEAM, further explores learning-based mutation selection (Tian et al., [n. d.]). However, these approaches suffer from the challenges in generating syntactically correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al., who proposed BERT (Degiovanni and Papadakis, 2022). BERT generates mutation by masking the code elements to be mutated and then predicting the mutation via BERT (Devlin et al., 2018).
@@
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently investigated the use of LLMs for mutation generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutations. In contrast, our study conducts a broader comparison by evaluating LLMs against a diverse set of baselines, including both rule-based and learning-based approaches. Furthermore, we explore different prompt strategies, such as incorporating few-shot examples and enforcing structured JSON output, and we systematically examine multiple prompt templates in the context of mutation testing. Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al. (Dakhel et al., 2024) and Ibrahimzada et al. (Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. BugFarm (Ibrahimzada et al., 2023) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.

### v5

@@ 4.2. Effectiveness on High-Order Mutant Generation
Previously, we only evaluated LLMs in generating first-order mutants (FOMs). To explore the effectiveness of LLMs in generating high-order mutants (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating LLMut on the recent mainstream models, including GPT-4o and DS-671b.
We modified the original prompt to explicitly request HOM generation, shown as Figure 4. The prompts directly instruct LLMs to generate HOMs, and each HOM is composed of several first-order mutants.
@@
Table 19 shows the results. For the effectiveness metrics, HOMs generally achieve average Ochiai scores and higher coupling rates across these models, indicating closer semantic alignment with real bugs. However, FOMs achieve higher real bug detectability on all models. For the validity metrics, HOMs exhibit lower duplicate mutant rates, suggesting improved mutant diversity. However, FOMs exhibit significantly higher compilability rates, which aligns with the expectation that larger code changes are more prone to introducing compilation errors. Our preliminary experiments suggest that strengthening LLM-based HOM generation is promising, as HOMs are semantically closer to real bugs and face more challenges in maintaining syntactic correctness.
@@ 4.3. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutant generation approaches and suggest potential improvements.
@@ 4.3.1. Advantages of LLM-based Approaches
Our study demonstrates that LLMs can generate high-quality mutants for mutation testing at an acceptable cost. The key advantage of LLMs lies in their ability to produce mutants with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
Additionally, LLMs could generate syntactically diverse mutants compared with traditional approaches, which is potentially the reason for generating high-coupling mutants. For example, Listing 1 presents two compilable mutants generated by GPT for Chart-1, which are theoretically beyond the capabilities of traditional approaches. In Mutant-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly. This mutant is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement. Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutants. In Mutant-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutant generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
Moreover, our study demonstrates that LLMs have clear potential in generating HOMs. Compared with first-order mutants (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability. The results highlight that strengthening LLM-based HOM generation is therefore a promising research direction.
@@ 4.3.2. Advantages of Traditional Approaches
Despite the advantages of LLM-generated mutants, traditional approaches still exhibit strengths in reliability and efficiency. They produce fewer non-compilable and duplicate mutants, with PIT achieving a 100% compilability rate and a 0% duplicate mutant rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutant rate. In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutant rates. Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutant rate, further avoiding redundant costs. This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
@@ 4.3.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate that there is still substantial room for improvement. One major limitation is the high rate of non-compilable mutants, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutant generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutants before execution can enhance their reliability.
Moreover, exploring the applications of LLM-generated mutants in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL). Our study has shown that LLM-generated mutants often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
Another important direction is cross-language generalization. While our study primarily focuses on Java, mutation testing is equally critical in languages such as C/C++, Python, JavaScript, and Rust. Traditional mutation testing tools usually require language-specific implementations, leading to duplicated engineering efforts and limited portability. In contrast, modern LLMs possess inherent multilingual capabilities, as they are pre-trained on large-scale, cross-language code corpora. This enables them to naturally transfer mutation strategies across programming languages without retraining from scratch. Future work should investigate how well LLMs can exploit this cross-language knowledge, and how prompts, training data, or filtering strategies can be adapted to different language paradigms (e.g., strongly typed vs. dynamically typed). Advancing towards multilingual mutant generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
@@ 4.4. Limitations of Our Study
First, while we evaluated a comprehensive set of static and behavioral metrics, we did not investigate more advanced dynamic indicators. For instance, the phenomenon of subsuming mutants (Garg et al., 2022; Diniz et al., 2021), where one mutant is always killed by any test that kills another, has been studied in the mutant testing literature as a way to assess better redundancy and the true fault-detection power of test suites. Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutants beyond the metrics reported in this paper. Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025a), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
Second, our experiments are limited to the Java programming language. Although Java is the most widely studied language in mutation testing, it does not cover the diversity of programming paradigms. Mutant behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments. Future research should examine the generalizability of LLM-based mutant generation across multiple languages. Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutant generation a promising direction for broadening the applicability of our findings.
Third, our study relies on a fixed few-shot examples, which restricts LLMs to producing context-specific mutants for the given code snippets. Moreover, constructing few-shot examples requires manual effort and may introduce unintended biases into the model’s behavior during inference. In addition, we impose no restrictions on the mutation operators applied by LLMs, whereas rule-based approaches rely on predefined and relatively simple operators. Consequently, it is still an open question how LLMs would perform if constrained to the same operator set as rule-based approaches.
@@ 5. Threats to Validity
The selected programming language, mutant generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT models, StarCoder, CodeLlama, and DeepSeek models). Moreover, our study further covers two state-of-the-art LLM-based mutant generation approaches, BugFarm and LLMorpheus, as baselines, both of which are equipped with the recent mainstream models GPT-4o and DS-671b.
@@
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014b) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset, ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of most LLMs we use (except DeepSeek-V3-671b only) and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutant that is an exact match with the faults we investigate. On Defects4J, the results are LLMut (GPT-3.5) (133), LLMut (GPT-4o) (63), LLMut (GPT-4o-M) (138), LLMut (DS-671b) (72), LLMut (DS-236b) (66), LLMut (SC-16b) (13), LLMut (CL-13b) (28), BugFarm (GPT-4o) (59), BugFarm (DS-671b) (66), LLMorpheus (GPT-4o) (26), LLMorpheus (DS-671b) (31), LEAM (287), Bert (12), and Major (22), respectively. On ConDefects, the results are LLMut (GPT-3.5) (22), LLMut (GPT-4o) (25), LLMut (GPT-4o-M) (31), LLMut (DS-671b) (35), LLMut (DS-236b) (32), LLMut (SC-16b) (35), LLMut (CL-13b) (56), BugFarm (GPT-4o) (22), BugFarm (DS-671b) (29), LLMorpheus (GPT-4o) (19), LLMorpheus (DS-671b) (28), LEAM (56), Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, LEAM tends to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with LLMut (CL-13b). Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for LLM-based approaches with the GPT models and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on both datasets, achieving a Spearman coefficient of 0.943 and a Pearson correlation of 0.944, both with a -value less than 0.05, indicating their performance is similar on both datasets.
@@
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompt design, context length, few-shot examples, and mutant numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs.
@@
The subjective nature of human decisions in labeling for equivalent mutants, non-compilation errors, and surviving mutants is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutant. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two authors.
@@ 6. Related Work
6.1. Mutation Testing and Mutant Generation
@@
Mutation testing is a systematic testing approach that aims to guide the generation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutants) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978), one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can support multiple applications in software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutants are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025a), test prioritization (Shin et al., 2019) and program repair (Hanna et al., 2025; Xiao et al., 2023, 2024).
@@
All these tasks require high-utility mutants. This is because it is required to compile and execute every mutant with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutants, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2016, 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutants are used as substitutes for real defects, the mutants must be as syntactically close to the real bugs as possible, making the mutants more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of mutants must be coupled with real faults/bugs to make these techniques effective.
@@
Traditional mutation testing approaches generate mutants by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), MIN (Zhang et al., 2025), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
@@
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutants. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. LEAM++, the extension of LEAM, further explores learning-based mutation selection (Tian et al., [n. d.]). However, these approaches suffer from challenges in generating syntactically correct mutants and mutants with diverse forms, often failing to produce syntactically valid mutants. Perhaps the closest to our work is that of Degiovanni et al., who proposed BERT (Degiovanni and Papadakis, 2022). BERT generates mutants by masking the code elements to be mutated and then predicting the mutants via BERT (Devlin et al., 2018).
@@
Several recent studies have also explored the use of LLMs for mutant generation. Tip et al. concurrently investigated the use of LLMs for mutant generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutants. In contrast, our study conducts a broader comparison by evaluating LLMs against a diverse set of baselines, including both rule-based and learning-based approaches. Furthermore, we explore different prompt strategies, such as incorporating few-shot examples and enforcing structured JSON output, and we systematically examine multiple prompt templates in the context of mutation testing. Similarly, Endres et al. employed large models for mutant generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al. (Dakhel et al., 2024) and Ibrahimzada et al. (Ibrahimzada et al., 2025) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. BugFarm (Ibrahimzada et al., 2025) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM. Compared with these approaches, our study is more comprehensive and evaluates existing mutant generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutants using established metrics, evaluate the syntactic diversity of mutants, and analyze the non-compilable mutants produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutant generation, and highlight directions for future improvement.